Using the method of substitution to help solve for this equation.
x+3y=1
-2x+2y=6
I am stuck this is what I have so far
x+3(-2x+6)=1
x-6+24=1
Am I on the right track?

Answer by Alan3354(69443)

Using the method of substitution to help solve for this equation.
x+3y=1
-2x+2y=6
I am stuck this is what I have so far
x+3(-2x+6)=1
x-6+24=1
Am I on the right track?
=================
You subbed (-2x+6) for y, that won't work.
---
-2x+2y=6
-x + y = 3
y = x+3
----
x+3y=1
x + 3(x+3) = 1
etc.